OpenStudy (anonymous):

what is the y - intercept of the graph of 4x + 2y = 12

OpenStudy (anonymous):

2y = -4x + 12 y = -2x + 6 you get the y intercept when x is zero so plug it in y = -2(0) + 6 y = 6 so it's (0, 6)

OpenStudy (amistre64):

6..i agree....6

OpenStudy (anonymous):

^_^